Compared to the focus and directrix of y = ax^2, explain how the focus and directrix of
(y-k) = a(x-h)^2 are affected by the values of h and k.

Answer by stanbon(75887)

Compared to the focus and directrix of y = ax^2, explain how the focus and directrix of
(y-k) = a(x-h)^2 are affected by the values of h and k.
-------------------
x^2 = ay
4p = a
p = a/4
Focus:: (0,a/4)
Directris: y = -a/4
===============================
y-k = a(x-h)^2
Focus:: (0+h,(a/4)+k)
Directrix: y = (-a/4)+k
==============================
Cheers,
Stan H.
==================